Previous

Calendar

September

2006

Next
Sunday Monday Tuesday Wednesday Thursday Friday Saturday

27

JD239

28

JD240
hailstorm in the afternoon (see photos on 060829)

29

JD241 (blue skies)
Site Visit (SpB, SS): Details , Photos (25)

from 11:04-11:20 MST, Changed profiler (li-6251) dessicant

Again modified TC23x logger program, now running "tc_05.csi".

from 15:30-16:30 MST, Replaced N2 at tower (New cylinder= CC153284, Old cylinder= CC199739)

Checked East Robot (done=0, battery/power is ok)

30

JD242

31

JD243
Site Visit (SpB, JH, MP): Details , Photos (11)

Again modified TC23x logger program, now running "tc_06.csi".

Swapped East Robot Flasks (done=1, note: recycled flasks from last week)

1

JD244

2

JD245

3

JD246
Site Visit (SpB, JB, LL, SO, A): Details , Photos I (4), , Photos II (23)

Moved around tubing for Hydra.

from 12:12-12:30 MST, Added Hydra/opto22 back to quacker (start ops29)

4

JD247
CU Holiday (Labor Day)

5

JD248
Site Visit (SpB, AT): Details , Photos (26)

Hydra: at ~14:00 MST, started collecting hydra data (using 5 inlets). [hydra_prog_como_test.txt]

from 15:00-15:30 MST, removed broken am25t (sn 1616) from tc23x logger

6

JD249
Hydra: at ~10:20 MST, modified hydra program [hydra_prog_como_Aonly.txt]

7

JD250
Site Visit (JH, MP, AT): Details

(Full "Corn" Moon)

8

JD251(wet/overcast)
Site Visit (SpB): Details , Photos (26)

Hydra: at ~10:30 MST, moved Como Creek inlets up to 1m (from 0.5m)

Swapped East Robot Flasks (done=1)

9

JD252
Site Visit (SpB, JH, MP, KS, JB, DM, JM, M, R): Details , Photos (16)

One Tree Harvested

10

JD253

11

JD254

12

JD255

13

JD256
Site Visit (SpB): Details , Photos (27)

Modified CNR logger program (saving more parameters), now running "rad_02.csi".

Hydra: connected more inlet lines

14

JD257
Site Visit (SpB, JH, MP): Details , Photos (26)

Hydra: increased ref gas (N2) flow rate

Adjusted location of li-7500 at 2.6m

Swapped East Robot Flasks (done=1)

15

JD258

16

JD259

17

JD260

18

JD261
Site Visit (SpB): Details , Photos (6)

from 12:20-12:32 MST, Changed profiler (li-6251) dessicant

turned off ozone flux system (a mistake!)

19

JD262
Site Visit (SpB): Details , Photos (11)

Modified UCB logger program (added Honeywell Pabs sensor), now running "ucb_06.csi".

Hydra: at 10:20 MST, dropped dry ice in hydra box

turned back on ozone flux system

20

JD263
Site Visit (SpB): Details , Photos (1)

removed broken am25t (sn 1396) from tc23x logger (ie, the one on the ground)

21

JD264

22

JD265 (snow!)
Fall Equinox

23

JD266
Site Visit (SpB, JB, KS, AS, T): Details , Photos (18)

from 10:15-11:15 MST, Replaced N2 at tower (New Cylinder= CC140777, Old cylinder= CC153284)

Hydra: shut down and moved to the North Canopy Tower.

East robot flasks swapped, and East Robot removed (Andy S.)

24

JD267
Site Visit (SpB): Details , Photos (17)

added repaired am25t (sn 1616) back to tc23x logger, now running "tc_07.csi"

25

JD268

26

JD269

27

JD270

28

JD271 (windy)
Site Visit (SpB): Details , Photos (5)

moved am25t (sn 1616) to outside of tc23x box.

29

JD272
Site Visit (SpB): Details , Photos (10)

connected all TCs to am25t (sn 1616). TC I_11 could not reach so that is now connected to tc23x, diff chan11, now running "tc_08.csi"

30

JD273